# Bridging Diffusion Models and 3D Representations: A 3D Consistent Super-Resolution Framework

### v1

@@ Abstract
We propose 3D Super Resolution (3DSR), a novel 3D Gaussian-splatting-based super-resolution framework that leverages off-the-shelf diffusion-based 2D super-resolution models. 3DSR encourages 3D consistency across views via the use of an explicit 3D Gaussian-splatting-based scene representation. This makes the proposed 3DSR different from prior work, such as image upsampling or the use of video super-resolution, which either don’t consider 3D consistency or aim to incorporate 3D consistency implicitly. Notably, our method enhances visual quality without additional fine-tuning, ensuring spatial coherence within the reconstructed scene. We evaluate 3DSR on MipNeRF360 and LLFF data, demonstrating that it produces high-resolution results that are visually compelling, while maintaining structural consistency in 3D reconstructions. Code will be released.
@@
Figure 2: Motivation: Diffusion-based super-resolution (SR) methods enhance details in high-resolution (HR) images but fail to maintain 3D consistency across views. Given a set of low-resolution (LR) images (a), we apply StableSR [32] to generate super-resolved images (b). However, the SR results introduce hallucinated details, as seen in the distorted steel strings of the bike, which appear misaligned and inconsistent. When rendered with 3DGS in (c), these inconsistencies lead to blurring and incorrect geometry. In the second row, the SR output (b) significantly deviates from the ground truth (d), demonstrating the instability of diffusion-based SR. Across multiple views, the inconsistencies in hallucinated textures further degrade 3DGS rendering, resulting in incorrect and incoherent textures.
@@ 1 Introduction
Recent advances in 3D representation learning have significantly improved novel view synthesis (NVS), enabling realistic rendering of scenes from sparse observations. Concretely, techniques such as Neural Radiance Fields (NeRFs) [24] and Gaussian Splatting (GS) [14] have demonstrated remarkable success in reconstructing high-fidelity 3D representations, benefiting applications in virtual reality, augmented reality, robotics, and content creation. However, despite their effectiveness, these methods are inherently limited by the spatial resolution of the input camera views, resulting in synthesized images that often lack fine details. This limitation restricts their applicability in high-quality rendering and immersive real-world experiences.
To address this resolution constraint, a straightforward approach is to upsample low-resolution (LR) training data using either image super-resolution (ISR) [18, 44, 17, 42, 33, 35, 37, 32] or video super-resolution (VSR) [34, 20, 7, 45, 39] techniques. However, applying ISR methods independently across views often introduces inconsistencies, leading to artifacts and a loss of geometric coherence. In contrast, VSR methods, while designed to enforce temporal consistency explicitly and 3D consistency implicitly, tend to struggle to preserve fine-grained details, resulting in blurred textures compared to ISR techniques. An alternative solution is to refine the underlying 3D representation directly, enhancing scene details at the feature level. However, existing approaches [27] often require significant computational resources or introduce new artifacts due to the inaccurate reconstruction of high-frequency details.
In parallel, diffusion models [11, 29, 30, 21, 19, 1, 2] have recently demonstrated remarkable results, often topping benchmarks in image synthesis, editing, and super-resolution. Unlike traditional deep learning models, diffusion-based approaches iteratively refine noisy inputs, enabling the generation of high-frequency textures and realistic details that surpass conventional GAN- or CNN-based super-resolution methods. Diffusion-based image super-resolution [37, 32] has significantly improved production of photo-realistic details with sharper textures. However, these fine details are often hallucinated rather than faithfully reconstructed, leading to cross-view inconsistency when applied to multi-view image generation. Since diffusion models lack an inherent understanding of 3D geometry, applying them independently to different viewpoints results in inconsistent textures and structural artifacts across synthesized views, akin to classic ISR.
@@
To evaluate our approach, we conduct extensive experiments on two real-world datasets, MipNeRF360 [6] and LLFF [23], and assess performance using PSNR, SSIM, LPIPS [43], NIQE [25], MEt3R [3], and FID [10]. As shown in Figure 1 and Figure 2, experimental results demonstrate that the proposed method improves perceptual quality, geometric consistency, and rendering fidelity compared to existing state-of-the-art techniques.
@@ 2 Related Work
Recent advances in 3D representations have significantly improved novel view synthesis (NVS), enabling high-quality scene rendering from sparse inputs. Neural Radiance Fields (NeRFs) [24] model scenes as continuous volumetric functions, achieving photo-realistic synthesis but requiring dense input views and long training times. Variants such as Mip-NeRF [5] and Instant-NGP [26] improve anti-aliasing and training efficiency, yet resolution constraints remain a challenge, leading to blurry or low-detail renderings at high resolutions.
3D Gaussian Splatting (3DGS) [14] offers a more efficient alternative, representing scenes as adaptive 3D Gaussians, enabling real-time rendering without ray marching. Mip-Splatting [41] further enhances 3DGS by integrating multi-scale representations, reducing aliasing and improving view extrapolation.
@@ 2.2 Super-Resolution
Image Super-Resolution (ISR). ISR aims to reconstruct a high-resolution (HR) image from a single LR input. Traditional CNN-based ISR methods, such as EDSR [18] and RCAN [44], leverage residual learning and channel attention to recover fine-grained textures. Transformer-based SR models, including SwinIR [17] and Restormer [42], have further advanced the field by utilizing self-attention mechanisms to improve feature aggregation and enhance reconstruction quality. Generative models, such as ESRGAN [33] and GFP-GAN [35], employ generative adversarial networks (GANs) to address the inherently ambiguous nature of the super-resolution problem. To synthesize photo-realistic details, GAN-based methods hallucinate fine textures and improve perceptual quality. However, these models are typically focused on specific categories, such as faces [35], and may struggle to generalize across more complex real-world SR tasks. More recently, diffusion models have emerged as a promising alternative, surpassing traditional discriminative and GAN-based methods in ISR. Instead of directly learning a mapping between LR and HR images, diffusion-based SR models iteratively refine images through a denoising process. Models like SinSR [37] and StableSR [32] excel in producing highly detailed and texture-rich images in large-scale real world datasets. However, a key limitation of diffusion-based ISR methods is their lack of geometric awareness, which can lead to challenges in maintaining structural consistency, particularly when applied to 3D scene synthesis.
@@
Video Super-Resolution (VSR). In VSR, the goal is to enhance video sequences while preserving temporal consistency. EDVR [34] and BasicVSR++ [7] improve spatio-temporal feature aggregation, effectively reducing flickering artifacts in enhanced frames. Recent transformer-based methods, such as TTVSR [20], integrate spatial and temporal self-attention to achieve strong performance in enhancing video resolution. Recently, VideoGigaGAN [39] has emerged as a promising method, employing a generative adversarial network (GAN) to upscale videos focusing on both perceptual quality and temporal consistency. This approach has demonstrated impressive results in preserving high-frequency details while maintaining coherence across multiple frames. Additionally, Upscale-A-Video [45], another recent method, adopts a text-based diffusion model and improves temporal coherence through local temporal layers and a global recurrent latent propagation module. Despite these advances, VSR methods often enforce temporal consistency via implicit spatio-temporal aggregation mechanisms [34, 20, 7, 45] or flow-guided propagation [39, 45], but they do not guarantee 3D view consistency. This lack of geometric awareness can lead to inconsistencies in spatial coherence when applied to multi-view scene reconstruction. Furthermore, VSR techniques frequently require multi-frame alignment, which increases computational complexity, particularly when handling high-resolution videos or large-scale 3D scene synthesis. These challenges can result in jittery, less detailed reconstructions and higher computational costs.
@@
Super-Resolution in 3D. Recent works have explored enhancing the quality of 3D assets using super-resolution techniques, but naively applying ISR methods to individual frames often leads to 3D inconsistencies, as these methods treat each view independently, disregarding geometric coherence. To maintain consistency with the LR input, some methods incorporate subpixel loss functions that align super-resolved outputs with the original LR images, such as NeRF-SR [31] and SRGS [9], which preserve 2D details but fail to ensure cross-view consistency, potentially causing structural distortions in novel views. To address multi-view consistency, some approaches leverage temporal coherence from VSR, like SuperGaussian [27], which extends VSR principles to 3D Gaussian Splatting but requires fine-tuning the video model, limiting its general applicability. It relies on frame-based SR techniques and is constrained by the temporal limitations of video models, which don’t explicitly address 3D scene consistency. Additionally, some methods enhance model representation when high-resolution input data is available, such as 4K-NeRF [38] and UHD-NeRF [16]. This improves rendering quality with high-resolution NeRF architectures. Further, SRGS [9] and GaussianSR [12, 40] focus on generating high-resolution details for 3D Gaussians. But these methods improve the 3D representation rather than ensuring 3D consistency, which distinguishes them from our goal of upsampling a 3D scene in a 3D-consistent manner.
@@ 2.3 View Consistent Diffusion Model
3D Object-Level Consistency. Several recent works have tackled 3D-consistent object synthesis using diffusion models. SyncDreamer [22] and MVDream [28] extend text-to-image diffusion models by enforcing multi-view consistency at the object level, enabling high-quality, consistent 3D object generation from a few input views. These methods typically leverage image-space priors and cross-view attention mechanisms to generate multiple perspectives of an object, while preserving structural consistency. However, they primarily focus on isolated objects rather than entire scenes. As a result, they struggle to model complex spatial relationships between multiple elements in a 3D environment, limiting their applicability to full-scene generation.
2D Multi-View Consistency. Other approaches emphasize enforcing consistency at the 2D level by refining multi-view diffusion outputs. MultiDiffusion [4] improves spatial consistency in image generation by synchronizing the denoising process across overlapping regions of multiple views. Generative Power of Ten [36] extends this concept by incorporating a joint multi-scale sampling approach, averaging noise across multiple views from different scales to encourage alignment. While these techniques improve 2D consistency, they cannot be directly applied to 3D scenes due to the lack of explicit geometry modeling, which results in missing feature correspondences across different views.
@@ 3 Method
We propose 3DSR, a 3D Gaussian-splatting-based super-resolution framework that leverages diffusion-based 2D super-resolution to enhance 3D consistency across different views. Note, the diffusion-based SR method used in this framework is not constrained to single image super-resolution (ISR) or video super-resolution (VSR), enabling 3DSR to benefit from future advances in this area. Compared to VSR-based methods like SuperGaussian [27], our approach encourages 3D consistency more explicitly without fine-tuning the video model.
@@ 3.3 Training Objective
To optimize 3D Gaussian Splatting (3DGS) for higher fidelity, we adopt the subsampling-based regularization strategy proposed in [31, 9]. This approach mitigates inconsistencies in high-resolution details by subsampling the rendered image $R^{i}$ to match the resolution of the low-resolution input $L^{i}$, producing $R_{\text{lr}}^{i}$. This encourages alignment between the rendered and the input images, improving overall consistency during rendering optimization.
@@
For both $\mathcal{L}_{\text{hr}}$ and $\mathcal{L}_{\text{lr}}$, we follow Mip-Splatting [41], employing a combination of an $\mathcal{L}_{1}$ loss and a D-SSIM loss $\mathcal{L}_{\text{D-SSIM}}$, i.e.,
@@ 4.1 Experimental Setup
Dataset. We evaluate our approach on two real-world datasets, MipNeRF360 [6] and LLFF [23]. The MipNeRF360 dataset consists of high-resolution images at its native $\times 1$ scale (approximately $3000\times 4000$ pixels) and includes both indoor and outdoor scenes. To simulate different resolution scenarios, we apply two downsampling factors: $\times 16$ with subsequent upsampling to $\times 4$, and $\times 8$ with upsampling to $\times 2$. This use of two downsampling factors encompasses two distinct scenarios: one in which the input consists of low-resolution images, and another where moderately high-resolution images are used for upsampling.
In contrast, the LLFF dataset contains images captured from forward-facing camera poses and provides significantly fewer frames for both training and testing (approximately 20204040) compared to MipNeRF360, which provides around 150150200200 frames. We adhere to the testing protocol of DiSR-NeRF [15], where images are downsampled by a factor of $\times 8$ and subsequently upsampled to $\times 2$.
To assess rendering quality, for both datasets we follow the protocol used by prior work [24, 14], removing every $8$th frame from the training set and designating it as a test image.
@@
Metrics. We evaluate our method using traditional metrics such as PSNR, SSIM, and LPIPS [43], the no-reference image quality metric NIQE [25], the 3D consistency metric MEt3R [3], and the FID [10], which measures the distributional similarity between generated and ground-truth images in feature space.
Baselines. We compare our approach with three representative baselines: StableSR [32], a diffusion-based single-image super-resolution (ISR) method; DiSR-NeRF [15], a diffusion-based 3D super-resolution method; and SuperGaussian [27], a video super-resolution (VSR)-based approach for 3D super-resolution.
StableSR serves as a baseline to evaluate 3D consistency in single-image-based SR. Since each image is super-resolved independently, it does not account for inter-view coherence, often leading to 3D inconsistencies. To evaluate StableSR, we super-resolve each low-resolution (LR) image independently and feed the resulting high-resolution (HR) images into Mip-Splatting [41] for 3D reconstruction.
@@
SuperGaussian, uses a VSR-based pipeline, leveraging video models to promote 3D consistency. It enhances a sequence of LR images via a video SR model before 3D reconstruction. Since our datasets are not temporally consistent, we synthesize a continuous video sequence by interpolating three intermediate frames between each pair of training views. We render a low-resolution video along this trajectory and apply a VSR model [8], then use its output frames to train a Gaussian Splatting model [13].
Implementation details. In our experiments, we adopt Mip-Splatting [41] as 3D representation and follow its original setup, training for 30,000 iterations using LR images. For our method, we use StableSR-Turbo [32], a diffusion-based SR model trained to produce high-quality outputs in just four denoising steps, significantly improving runtime efficiency without compromising quality. We set the diffusion sampling steps to 4, with 5,000 3DGS training iterations per step. During each step, we set the subsampling weight to 1, $\lambda=1$, and $\delta=0.2$. All experiments are conducted on an Nvidia A6000 GPU with 49GB of memory.

### v2

@@ Abstract
We propose 3D Super Resolution (3DSR), a novel 3D Gaussian-splatting-based super-resolution framework that leverages off-the-shelf diffusion-based 2D super-resolution models. 3DSR encourages 3D consistency across views via the use of an explicit 3D Gaussian-splatting-based scene representation. This makes the proposed 3DSR different from prior work, such as image upsampling or the use of video super-resolution, which either don’t consider 3D consistency or aim to incorporate 3D consistency implicitly. Notably, our method enhances visual quality without additional fine-tuning, ensuring spatial coherence within the reconstructed scene. We evaluate 3DSR on MipNeRF360 and LLFF data, demonstrating that it produces high-resolution results that are visually compelling, while maintaining structural consistency in 3D reconstructions. Project page: https://consistent3dsr.github.io/.
Figure 1: Qualitative comparison on the LLFF dataset with a downsampling factor of $\times 8$ and upsampling of $\times 4$. We train 3DGS with low-resolution (LR) inputs and with 3DSR (ours) and render test views. The results demonstrate that our method yields fewer artifacts from 3D inconsistencies and better preserves overall structural integrity compared to the baselines.
@@
Figure 2: Motivation: Diffusion-based super-resolution (SR) methods enhance details in high-resolution (HR) images but fail to maintain 3D consistency across views. Given a set of low-resolution (LR) images (a), we apply StableSR [wang2024exploiting] to generate super-resolved images (b). However, the SR results introduce hallucinated details, as seen in the distorted steel strings of the bike, which appear misaligned and inconsistent. When rendered with 3DGS in (c), these inconsistencies lead to blurring and incorrect geometry. In the second row, the SR output (b) significantly deviates from the ground truth (d), demonstrating the instability of diffusion-based SR. Across multiple views, the inconsistencies in hallucinated textures further degrade 3DGS rendering, resulting in incorrect and incoherent textures.
@@ 1 Introduction
Recent advances in 3D representation learning have significantly improved novel view synthesis (NVS), enabling realistic rendering of scenes from sparse observations. Concretely, techniques such as Neural Radiance Fields (NeRFs) [mildenhall2021nerf] and Gaussian Splatting (GS) [kerbl3Dgaussians] have demonstrated remarkable success in reconstructing high-fidelity 3D representations, benefiting applications in virtual reality, augmented reality, robotics, and content creation. However, despite their effectiveness, these methods are inherently limited by the spatial resolution of the input camera views, resulting in synthesized images that often lack fine details. This limitation restricts their applicability in high-quality rendering and immersive real-world experiences.
To address this resolution constraint, a straightforward approach is to upsample low-resolution (LR) training data using either image super-resolution (ISR) [lim2017enhanced, zhang2018image, liang2021swinir, zamir2022restormer, wang2018esrgan, wang2021towards, wang2024sinsr, wang2024exploiting] or video super-resolution (VSR) [wang2019edvr, liu2022learning, chan2022basicvsr++, zhou2024upscaleavideo, xu2024videogigagan] techniques. However, applying ISR methods independently across views often introduces inconsistencies, leading to artifacts and a loss of geometric coherence. In contrast, VSR methods, while designed to enforce temporal consistency explicitly and 3D consistency implicitly, tend to struggle to preserve fine-grained details, resulting in blurred textures compared to ISR techniques. An alternative solution is to refine the underlying 3D representation directly, enhancing scene details at the feature level. However, existing approaches [shen2024supergaussian] often require significant computational resources or introduce new artifacts due to the inaccurate reconstruction of high-frequency details.
In parallel, diffusion models [ho2020denoising, song2021denoising, SongICLR2021, liu2023flow, LipmanICLR2023, albergo2023building, albergo2023stochastic] have recently demonstrated remarkable results, often topping benchmarks in image synthesis, editing, and super-resolution. Unlike traditional deep learning models, diffusion-based approaches iteratively refine noisy inputs, enabling the generation of high-frequency textures and realistic details that surpass conventional GAN- or CNN-based super-resolution methods. Diffusion-based image super-resolution [wang2024sinsr, wang2024exploiting] has significantly improved production of photo-realistic details with sharper textures. However, these fine details are often hallucinated rather than faithfully reconstructed, leading to cross-view inconsistency when applied to multi-view image generation. Since diffusion models lack an inherent understanding of 3D geometry, applying them independently to different viewpoints results in inconsistent textures and structural artifacts across synthesized views, akin to classic ISR.
@@
To evaluate our approach, we conduct extensive experiments on two real-world datasets, MipNeRF360 [barron2022mip] and LLFF [mildenhall2019llff], and assess performance using PSNR, SSIM, LPIPS [zhang2018unreasonable], NIQE [mittal2012making], MEt3R [asim2025met3r], and FID [heusel2017gans]. As shown in Figure 1 and Figure 2, experimental results demonstrate that the proposed method improves perceptual quality, geometric consistency, and rendering fidelity compared to existing state-of-the-art techniques.
@@ 2 Related Work
Recent advances in 3D representations have significantly improved novel view synthesis (NVS), enabling high-quality scene rendering from sparse inputs. Neural Radiance Fields (NeRFs) [mildenhall2021nerf] model scenes as continuous volumetric functions, achieving photo-realistic synthesis but requiring dense input views and long training times. Variants such as Mip-NeRF [barron2021mip] and Instant-NGP [mueller2022instant] improve anti-aliasing and training efficiency, yet resolution constraints remain a challenge, leading to blurry or low-detail renderings at high resolutions.
3D Gaussian Splatting (3DGS) [kerbl3Dgaussians] offers a more efficient alternative, representing scenes as adaptive 3D Gaussians, enabling real-time rendering without ray marching. Mip-Splatting [yu2024mip] further enhances 3DGS by integrating multi-scale representations, reducing aliasing and improving view extrapolation.
@@ 2.2 Super-Resolution
Image Super-Resolution (ISR). ISR aims to reconstruct a high-resolution (HR) image from a single LR input. Traditional CNN-based ISR methods, such as EDSR [lim2017enhanced] and RCAN [zhang2018image], leverage residual learning and channel attention to recover fine-grained textures. Transformer-based SR models, including SwinIR [liang2021swinir] and Restormer [zamir2022restormer], have further advanced the field by utilizing self-attention mechanisms to improve feature aggregation and enhance reconstruction quality. Generative models, such as ESRGAN [wang2018esrgan] and GFP-GAN [wang2021towards], employ generative adversarial networks (GANs) to address the inherently ambiguous nature of the super-resolution problem. To synthesize photo-realistic details, GAN-based methods hallucinate fine textures and improve perceptual quality. However, these models are typically focused on specific categories, such as faces [wang2021towards], and may struggle to generalize across more complex real-world SR tasks. More recently, diffusion models have emerged as a promising alternative, surpassing traditional discriminative and GAN-based methods in ISR. Instead of directly learning a mapping between LR and HR images, diffusion-based SR models iteratively refine images through a denoising process. Models like SinSR [wang2024sinsr] and StableSR [wang2024exploiting] excel in producing highly detailed and texture-rich images in large-scale real world datasets. However, a key limitation of diffusion-based ISR methods is their lack of geometric awareness, which can lead to challenges in maintaining structural consistency, particularly when applied to 3D scene synthesis.
@@
Video Super-Resolution (VSR). In VSR, the goal is to enhance video sequences while preserving temporal consistency. EDVR [wang2019edvr] and BasicVSR++ [chan2022basicvsr++] improve spatio-temporal feature aggregation, effectively reducing flickering artifacts in enhanced frames. Recent transformer-based methods, such as TTVSR [liu2022learning], integrate spatial and temporal self-attention to achieve strong performance in enhancing video resolution. Recently, VideoGigaGAN [xu2024videogigagan] has emerged as a promising method, employing a generative adversarial network (GAN) to upscale videos focusing on both perceptual quality and temporal consistency. This approach has demonstrated impressive results in preserving high-frequency details while maintaining coherence across multiple frames. Additionally, Upscale-A-Video [zhou2024upscaleavideo], another recent method, adopts a text-based diffusion model and improves temporal coherence through local temporal layers and a global recurrent latent propagation module. Despite these advances, VSR methods often enforce temporal consistency via implicit spatio-temporal aggregation mechanisms [wang2019edvr, liu2022learning, chan2022basicvsr++, zhou2024upscaleavideo] or flow-guided propagation [xu2024videogigagan, zhou2024upscaleavideo], but they do not guarantee 3D view consistency. This lack of geometric awareness can lead to inconsistencies in spatial coherence when applied to multi-view scene reconstruction. Furthermore, VSR techniques frequently require multi-frame alignment, which increases computational complexity, particularly when handling high-resolution videos or large-scale 3D scene synthesis. These challenges can result in jittery, less detailed reconstructions and higher computational costs.
@@
Super-Resolution in 3D. Recent works have explored enhancing the quality of 3D assets using super-resolution techniques, but naively applying ISR methods to individual frames often leads to 3D inconsistencies, as these methods treat each view independently, disregarding geometric coherence. To maintain consistency with the LR input, some methods incorporate subpixel loss functions that align super-resolved outputs with the original LR images, such as NeRF-SR [wang2022nerf] and SRGS [feng2024srgs], which preserve 2D details but fail to ensure cross-view consistency, potentially causing structural distortions in novel views. To address multi-view consistency, some approaches leverage temporal coherence from VSR, like SuperGaussian [shen2024supergaussian], which extends VSR principles to 3D Gaussian Splatting but requires fine-tuning the video model, limiting its general applicability. It relies on frame-based SR techniques and is constrained by the temporal limitations of video models, which don’t explicitly address 3D scene consistency. Additionally, some methods enhance model representation when high-resolution input data is available, such as 4K-NeRF [wang20224k] and UHD-NeRF [li2023uhdnerf]. This improves rendering quality with high-resolution NeRF architectures. Further, SRGS [feng2024srgs] and GaussianSR [hu2024gaussiansr, yu2024gaussiansr] focus on generating high-resolution details for 3D Gaussians. But these methods improve the 3D representation rather than ensuring 3D consistency, which distinguishes them from our goal of upsampling a 3D scene in a 3D-consistent manner.
@@ 2.3 View Consistent Diffusion Model
3D Object-Level Consistency. Several recent works have tackled 3D-consistent object synthesis using diffusion models. SyncDreamer [liu2023syncdreamer] and MVDream [shi2023mvdream] extend text-to-image diffusion models by enforcing multi-view consistency at the object level, enabling high-quality, consistent 3D object generation from a few input views. These methods typically leverage image-space priors and cross-view attention mechanisms to generate multiple perspectives of an object, while preserving structural consistency. However, they primarily focus on isolated objects rather than entire scenes. As a result, they struggle to model complex spatial relationships between multiple elements in a 3D environment, limiting their applicability to full-scene generation.
2D Multi-View Consistency. Other approaches emphasize enforcing consistency at the 2D level by refining multi-view diffusion outputs. MultiDiffusion [bar2023multidiffusion] improves spatial consistency in image generation by synchronizing the denoising process across overlapping regions of multiple views. Generative Power of Ten [wang2024generative] extends this concept by incorporating a joint multi-scale sampling approach, averaging noise across multiple views from different scales to encourage alignment. While these techniques improve 2D consistency, they cannot be directly applied to 3D scenes due to the lack of explicit geometry modeling, which results in missing feature correspondences across different views.
@@ 3 Method
We propose 3DSR, a 3D Gaussian-splatting-based super-resolution framework that leverages diffusion-based 2D super-resolution to enhance 3D consistency across different views. Note, the diffusion-based SR method used in this framework is not constrained to single image super-resolution (ISR) or video super-resolution (VSR), enabling 3DSR to benefit from future advances in this area. Compared to VSR-based methods like SuperGaussian [shen2024supergaussian], our approach encourages 3D consistency more explicitly without fine-tuning the video model.
@@ 3.3 Training Objective
To optimize 3D Gaussian Splatting (3DGS) for higher fidelity, we adopt the subsampling-based regularization strategy proposed in [wang2022nerf, feng2024srgs]. This approach mitigates inconsistencies in high-resolution details by subsampling the rendered image $R^{i}$ to match the resolution of the low-resolution input $L^{i}$, producing $R_{\text{lr}}^{i}$. This encourages alignment between the rendered and the input images, improving overall consistency during rendering optimization.
@@
For both $\mathcal{L}_{\text{hr}}$ and $\mathcal{L}_{\text{lr}}$, we follow Mip-Splatting [yu2024mip], employing a combination of an $\mathcal{L}_{1}$ loss and a D-SSIM loss $\mathcal{L}_{\text{D-SSIM}}$, i.e.,
@@ 4.1 Experimental Setup
Dataset. We evaluate our approach on two real-world datasets, MipNeRF360 [barron2022mip] and LLFF [mildenhall2019llff]. The MipNeRF360 dataset consists of high-resolution images at its native $\times 1$ scale (approximately $3000\times 4000$ pixels) and includes both indoor and outdoor scenes. To simulate different resolution scenarios, we apply two downsampling factors: $\times 16$ with subsequent upsampling to $\times 4$, and $\times 8$ with upsampling to $\times 2$. This use of two downsampling factors encompasses two distinct scenarios: one in which the input consists of low-resolution images, and another where moderately high-resolution images are used for upsampling.
In contrast, the LLFF dataset contains images captured from forward-facing camera poses and provides significantly fewer frames for both training and testing (approximately 20204040) compared to MipNeRF360, which provides around 150150200200 frames. We adhere to the testing protocol of DiSR-NeRF [lee2024disr], where images are downsampled by a factor of $\times 8$ and subsequently upsampled to $\times 2$.
To assess rendering quality, for both datasets we follow the protocol used by prior work [mildenhall2021nerf, kerbl3Dgaussians], removing every $8$th frame from the training set and designating it as a test image.
@@
Metrics. We evaluate our method using traditional metrics such as PSNR, SSIM, and LPIPS [zhang2018unreasonable], the no-reference image quality metric NIQE [mittal2012making], the 3D consistency metric MEt3R [asim2025met3r], and the FID [heusel2017gans], which measures the distributional similarity between generated and ground-truth images in feature space.
Baselines. We compare our approach with three representative baselines: StableSR [wang2024exploiting], a diffusion-based single-image super-resolution (ISR) method; DiSR-NeRF [lee2024disr], a diffusion-based 3D super-resolution method; and SuperGaussian [shen2024supergaussian], a video super-resolution (VSR)-based approach for 3D super-resolution.
StableSR serves as a baseline to evaluate 3D consistency in single-image-based SR. Since each image is super-resolved independently, it does not account for inter-view coherence, often leading to 3D inconsistencies. To evaluate StableSR, we super-resolve each low-resolution (LR) image independently and feed the resulting high-resolution (HR) images into Mip-Splatting [yu2024mip] for 3D reconstruction.
@@
SuperGaussian, uses a VSR-based pipeline, leveraging video models to promote 3D consistency. It enhances a sequence of LR images via a video SR model before 3D reconstruction. Since our datasets are not temporally consistent, we synthesize a continuous video sequence by interpolating three intermediate frames between each pair of training views. We render a low-resolution video along this trajectory and apply a VSR model [chan2022investigating], then use its output frames to train a Gaussian Splatting model [kerbl20233d].
Implementation details. In our experiments, we adopt Mip-Splatting [yu2024mip] as 3D representation and follow its original setup, training for 30,000 iterations using LR images. For our method, we use StableSR-Turbo [wang2024exploiting], a diffusion-based SR model trained to produce high-quality outputs in just four denoising steps, significantly improving runtime efficiency without compromising quality. We set the diffusion sampling steps to 4, with 5,000 3DGS training iterations per step. During each step, we set the subsampling weight to 1, $\lambda=1$, and $\delta=0.2$. All experiments are conducted on an Nvidia A6000 GPU with 49GB of memory.